There is a number of $2 and $5 coins. It is known that 1/3 of the number of $2 coins is greater than that of $5 coins by 1. If the total value of the coins is $39, (a) find the number of $2 coins. (b)find the total value of $5 coins
Answer by mananth(16949)

$2 coins = x
$5 coins y
x/3=y+1
x=3(y+1)
x=3y+3
2x+5y=39
1 x -3 y = 3 .............1
2 x + 5 y = 39 .............2
Eliminate y
multiply (1)by 5
Multiply (2) by 3
5 x -15 y = 15
6 x 15 y = 117
Add the two equations
11 x = 132
/ 11
x = 12.00 12
plug value of x in (1)
1 x + -3 y = 3
12 + -3 y = 3
-3 y = 3 -12
-3 y = -9
y = 3.00
$ 2 coins 12 numbers
$5 coins 3 numbers
m.ananth@hotmail.ca